OpenStudy (anonymous):

What is the value of y in the equation 6.4x + 2.8y = 44.4, when x = 3?

OpenStudy (anonymous):

First plug in '3' for 'x': 6.4(3) + 2.8y = 44.4 What's 6.4 * 3?

OpenStudy (anonymous):

19.2

OpenStudy (anonymous):

Yes, now we subtract it to both sides, what's 44.4 - 19.2?

OpenStudy (ribhu):

6.4*3+2.8y=44.4 (putting x =3 in the given equation) 19.2+2.8y=44.4 2.8y=44.4-19.2 2.8y=25.2 y=25.2/2.8 y=9

OpenStudy (anonymous):

-.- I thought you were the asker.

OpenStudy (anonymous):

What's 44.4 - 19.2? @Rafyy

OpenStudy (anonymous):

25.2

OpenStudy (anonymous):

Yes, so we have: 2.8x = 25.2 Now we divide 2.8 to both sides, what's 25.2 / 2.8?

OpenStudy (anonymous):

2.8y*

OpenStudy (anonymous):

9! =D

OpenStudy (anonymous):

Yep!